OpenStudy (anonymous):

What is the equation of the line passing through (–3, –9) and parallel to the line y = x + 6 in standard form? Answer x + y = 6 x – y = 6 x + y = –6 x – y = –6

OpenStudy (aroub):

parallel=same slope y=mx+b , m=1 , (-3,-9) -9=1(-3)+b -9=-3+b -6=b so the equation is y=x-6

OpenStudy (anonymous):

which one?

OpenStudy (aroub):

x-y=6